OpenStudy (anonymous):

Select the coordinates of two points on the line y = 8. (8, 8) and (8, 0) (8, -8) and (8, 8) (0, 8) and (8, 0) (0, 8) and (8, 8)

OpenStudy (anonymous):

is there an x coordinate as well?

OpenStudy (anonymous):

the general equation of line is y= mx + b where m is slope & b is constant Here we have y = 8 that is here slope m is zero we find slope by m = y2- y1 /x2 - x1 just check the points so m = 0 so here if we take (0, 8) and (8, 8) so 8 - 8 / 8 - 0 = 0

OpenStudy (anonymous):

@lauren_bisko naw that's all they give

OpenStudy (anonymous):

@tgawade is this the answer?

OpenStudy (anonymous):

Please check 1/0 = infinity ( not defined ) Yes that is answer

OpenStudy (anonymous):

ohhhhhhhh i get it now

OpenStudy (anonymous):

nvm i just needed to read it over again